# Traffic Flow and Speed Monitoring Based On Optical Fiber Distributed Acoustic Sensor

### v1

@@ Abstract
In the realm of intelligent transportation systems, accurate and reliable traffic monitoring is crucial. Traditional devices, such as cameras and lidars, face limitations in adverse weather conditions and complex traffic scenarios, prompting the need for more resilient technologies. This paper presents traffic flow monitoring method using optical fiber-based Distributed Acoustic Sensors (DAS). An innovative vehicle trajectory extraction algorithm is proposed to derive traffic flow statistics. In the processing of optical fiber waterfall diagrams, Butterworth low-pass filter and peaks location search method are employed to determine the entry position of vehicles. Subsequently, line-by-line matching algorithm is proposed to effectively track the trajectories. Experiments were conducted under both real highway and tunnel scenarios, showing that our approach not only extracts vehicle trajectories more accurately than the classical Hough and Radon transform-based methods, but also facilitates the calculation of traffic flow information using the low-cost acoustic sensors. It provides a new reliable means for traffic flow monitoring which can be integrated with existing methods.
Index Terms:
Optical Fiber Distributed Acoustic Sensor (DAS), Vehicle Detection, Trajectory Tracking, Vehicle Speed Estimation, Traffic Flow.
publicationid: pubid:
I Introduction
@@
Figure 1: Comparison of camera and optical fiber observations. Cameras cannot deal well with these situations but fiber systems are not affected.
@@
Nowadays, methods for monitoring traffic flow have predominantly turned to technologies such as cameras, radars, and lidars [5]. Deep Learning algorithms, including YOLO (2D multi-object detection) [6] and CenterPoint (3D multi-object detection) [7], are increasingly utilized for vehicle detection. Additionally, an effective combination of the Kalman filter and the Hungarian algorithm can be harnessed for state estimation and data association, thereby facilitating vehicle tracking and real-time positioning [8]. Nonetheless, as shown in Fig. 1.(a), the efficacy of these devices can be severely compromised by environmental conditions and their deployment strategies. Elements like rain, snow, fog, storms, and suboptimal lighting can distort or damage equipment, making round-the-clock, all-weather vehicle detection challenging. Compounding these issues, cameras, given their fixed viewing angles, introduce dead zones and issues relating to object sizes [9]. These problems are accentuated when detecting smaller objects nearby or larger objects at extended distances, as illustrated in Fig 1.(b). Present 2D detection methods often fall short in identifying distant small objects. While lidars can discern distance information over longer ranges, their high costs hinder widespread deployment, especially at highway ends [10]. Contrasting with cameras and lidars, which serve as “visual detectors”, underground optical fibers can be likened to “acoustic detectors”. Benefiting from numerous design advantages such as resistance to electromagnetic interference, heightened sensitivity, cost-effectiveness, durability, stealth, and adaptability, optical fibers may provide another reliable way for continuous and comprehensive traffic state perception in smart highways [11] and can be integrated with camera and Lidar-based methods.
In this study, we investigate methods using underground optical fiber-based Distributed Acoustic Sensing (DAS) for monitoring highway traffic, including highway and tunnel scenarios. The method mainly includes two parts: (1) vehicle detection and tracking by signal processing; and (2) traffic density, flow, and velocity estimation.
@@
In the processing of optical fiber data, the raw data is presented as three-dimensional waterfall diagrams. Then image preprocessing through min-max normalization and wavelet threshold denoising is presented. Based on these preprocessed optical fiber waterfall diagrams, we propose a vehicle position detection algorithm based on the Butterworth low-pass filter and peaks location search. Then vehicle trajectories are extracted based on line-by-line matching and polynomial fitting, thus achieving real-time detection and overall tracking for vehicles. We then present the calculation methods for three traffic flow information (i.e. velocity, flow and density) on road profile and road segment respectively, realizing traffic flow and speed monitoring on highways.
@@
The performances of the proposed methods are verified in real datasets collected from highway and tunnel scenarios. The results show that the vehicle tracking results extracted by our methods are more accurate than the traditional Hough and Radon transform-based methods. In the tunnel scenarios, our algorithm accuracy is up to 90.90%, which validates the usability of the proposed methods in practical scenarios considering the very low-cost signal processing based method. In complex vehicle conditions like congestion and illegal overtaking, the proposed methods can track vehicles when line overlapping and intersecting are present, validating the effectiveness of the proposed methods for complicated vehicle scenarios.
The main contributions of the paper are as follows:
@@
    We introduce the DAS system for vehicle detection, designed to gather data from both highways and tunnels.
@@
    We present a novel approach for vehicle detection and tracking, relying solely on optical fiber signals. This method exhibits remarkable accuracy, particularly within tunnel settings.
@@
    Our proposed technique effectively addresses challenges such as vehicle overtaking and congestion-induced crossover situations in tunnels.
II Related Works
Since distributed fiber optical sensing systems have been used in transportation [12], there have been some works using the DAS system for traffic monitoring, but these are mostly based on image and signal processing algorithms, such as Hough transform [13] and Radon transform [14]. Subsequent scholars are more committed to studying how to denoise the collected fiber signal, rather than analyzing the information in the fiber signal. In recent years, Deep Learning has made remarkable achievements in image denoising. An attention-based Convolutional Neural Network (CNN) structure has been introduced specifically for denoising short-time DAS signals [15]. However, these algorithms can only be used for denoising or line extracting, rather than monitoring changes in vehicle positions and real-time tracking. Additionally, their emphasis does not extend to the statistical analysis of traffic flow information. Moreover, DAS systems are commonly deployed in smart cities [16] and on campuses [17], where vehicles typically pass in sequence, and the situation is relatively straightforward. This configuration does not consider the challenges posed by long-distance installations on highways and the continuous tracking of vehicles over extended distances.
Most of the existing optical fiber signal processing relies on image processing techniques, such as the Hough transform and Radon transform, rather than processing at the pixel level. Because of the extensive radiation range of optical fiber signals, the Hough transform often identifies a single vehicle trajectory line as multiple lines. Conversely, the Radon transforms may overlook detecting lines with short driving distances. In addition, both the Hough transform and the Radon transform are primarily applicable for the detection of straight lines, making it challenging to discern instances of traffic congestion and speed variation.
In this paper, we propose a pixel-level approach to extract vehicle trajectories. Our designed method offers enhanced accuracy throughout the tracking process, effectively addressing challenges like crossovers and overlaps. Furthermore, we provide a comparative analysis of our proposed method with the Hough transform and Radon transform in Section V.
III Method
III-A Problem Formulation
@@
The original optical fiber data recorded by the DAS system can be represented as a matrix denoted by A:
@@
Fiber waterfall diagram: Three-dimensional waterfall diagrams after the visualization of the original optical fiber matrix A in Eq. (1) with the distance as the horizontal axis and time as the vertical axis.
@@
Fiber stripe: Clear banded stripes visible to the human eye in a fiber waterfall diagram. In this paper, fiber stripes represent the driving trajectories of the vehicles.
According to the above definition, we can draw optical fiber waterfall diagrams based on the collected fiber signals. In this section, we provide a visual representation of optical fiber waterfall diagrams in highway tunnel datasets, as shown in Fig. 2. In Fig. 2, the brightly colored lines, denoted as fiber stripes in the optical fiber waterfall diagrams, signify the spatiotemporal positions of vehicles. Highways, particularly in tunnel scenarios, often present unique conditions, such as congestion(the merging lines depicted in Fig. 2 (a) and (b)) and overtaking(the crossing lines depicted in Fig. 2 (c)). Accurately distinguishing the trajectory of each vehicle and conducting traffic flow statistics is a challenge that requires exploration. This is the primary objective of this paper.
Figure 2: Optical fiber waterfall diagrams illustrating vehicle congestion in a 2000-meter section of a highway tunnel dataset: (a) five vehicles congested in a jam, (b) six vehicles congested in a jam, and (c) one car overtaking illegally.
III-B Signal Preprocessing
The optical fiber data distribution among different datasets may vary differently, so signal preprocessing is needed to eliminate the background noise as much as possible to obtain a cleaner signal. In this section, we employ two techniques for signal preprocessing: min-max normalization and weighted wavelet threshold denoising.
III-B1 Normalization
Min-max normalization is one of the most popular techniques of data normalization, which is a linear transformation of the original data, mapping the original value x into the value in the interval [0, 1]. The formula is as follows:
${}x_{\rm{norm}}=\frac{x-\rm{min}}{\rm{max}-\rm{min}}$ (2)
where $\rm{max}$ and $\rm{min}$ are the maximum and minimum values of the sample data respectively.
The standardization method in this paper is to process the whole optical fiber waterfall diagram, rather than a single row or single column; otherwise, it will affect the subsequent discrimination of trucks and cars, scaling the matrix corresponding to the whole diagram to [0,1].
III-B2 Signal Denoising
@@
In signal processing, noise reduction is the primary task to be solved, and in our case, we employ a wavelet threshold denoising algorithm combining hard and soft threshold functions.
Due to its excellent time-frequency characteristics, wavelet theory has gained widespread use in signal processing and has experienced rapid development [18]. The energy of the signal corresponds to the wavelet coefficient with a larger amplitude, while the energy of noise corresponds to the wavelet coefficient with a smaller amplitude. Hence, an optimal threshold can be established, where wavelet coefficients exceeding this threshold are identified as valuable signals and are either retained or contracted. Conversely, wavelet coefficients falling below the threshold are deemed noise and are subsequently removed. Ultimately, the signal undergoes reconstruction through the inverse wavelet transform, resulting in the acquisition of a denoised signal.
For the selection of wavelet threshold function, a wavelet threshold denoising algorithm combining hard and soft threshold functions is employed to denoise vehicle vibration signals. The formula is as follows:
$\overline{{\omega}_{j,k}}=\begin{cases}{\omega}_{j,k}-a\lambda,&{\omega}_{j,k}\geq\lambda;\\ 0,&\lvert{\omega}_{j,k}\rvert\textless\lambda;\\ {\omega}_{j,k}+a\lambda,&{\omega}_{j,k}\leq-\lambda.\end{cases}$ (3)
@@
where ${\omega}_{j,k}$ is the wavelet coefficient, and $\lambda$ is the threshold. When $a=0$, it is a hard threshold function; when $a=1$, it is a soft threshold function; when $0\textless a\textless 1$, it represents a compromise between the hard and soft thresholds. In this paper, according to the actual situation, an ideal denoised diagram can be obtained when $a=0.5$.
The normalized and denoised results of different wavelet threshold denoising algorithms are shown in Fig. 3. After denoising, both the peak value and fundamental trend of the vibration signal are effectively retained.
Figure 3: Comparisions among different wavelet threshold denoising algorithms. (a) normalized diagram, (b) a = 0, (c) a = 1, and (d) a = 0.5.
III-C Vehicle Position Detection
While the preprocessing steps can effectively eliminate global noise and interference, achieving accurate extraction of structured information, including vehicle positions from the optical fiber stripe, requires additional filtering. Therefore, we employ a real-time vehicle positioning approach through a peak location search algorithm.
III-C1 Smoothing method of Butterworth Filter
The Butterworth filter, named after the British engineer and physicist Stephen Butterworth [19], is characterized by a frequency response curve that is flat to the maximum in the passband (i.e., a bandpass filter) without a ripple, gradually rolling off to zero in the stopband. The Butterworth low-pass filter can be expressed by the following formula for amplitude squared relative to frequency:
$\lvert H(\omega)\rvert^{2}=\frac{1}{1+\epsilon^{2}(\frac{\omega}{\omega_{c}})^{2n}}$ (4)
where n is the order of the filter, $\omega_{c}$ is the cutoff frequency, and $\omega$ is the operating frequency (passband frequency) of circuit, $\epsilon$ is the maximum passband gain. The value of $\epsilon$ is determined by the following formula:
${}H_{1}=\frac{H_{0}}{\sqrt{1+\epsilon^{2}}}$ (5)
where $H_{0}$ and $H_{1}$ represent the maximum and minimum passband gains, respectively.
It’s worth noting that the higher the order of the Butterworth filter, the steeper the amplitude-frequency response, resulting in a more effective filtering effect.
III-C2 Peaks location search algorithm
According to the features of the wave peaks, satisfying the principle that the first derivative is zero and the second derivative is negative, the position of the wave peaks can be searched as follows:
(1) Suppose the projection curve can be expressed as $V=[v_{1},v_{2},\ldots,v_{n}]$.
(2) Calculate the first-order difference vector ${\rm{Diff}}_{V}$:
${\rm{Diff}}_{v}(i)=V(i+1)-V(i)$ (6)
where $i\in 1,2,\ldots,N-1$.
(3) Apply the sign function to the difference vector:
${\rm{Trend}}={\rm{sign}}({\rm{Diff}}_{v})$ (7)
where
(4) Iterate through the $\rm{Trend}$ vector from the tail to perform the following:
${\rm{Trend}}(i)=\begin{cases}1,\text{if }{\rm{Trend}}(i)=0\text{ and }{\rm{Trend}}(i+1)\geq 0;\\ -1,\text{if }{\rm{Trend}}(i)=0\text{ and }{\rm{Trend}}(i+1)<0.\end{cases}$ (9)
(5) Perform the first-order difference operation on the $\rm{Trend}$ vector, as in Step 2, to obtain $R=\rm{Diff}(\rm{Trend})$.
(6) Traverse the difference vector $R$. If $R(i)=-2$, then $i+1$ is a crest of the projection vector $V$, and the corresponding crest is $V(i+1)$. If $R(i)=2$, then $i+1$ is a trough of the projection vector $V$, and the corresponding trough is $V(i+1)$.
After applying the Butterworth low-pass filter to the waterfall figure column-wise and subsequently extracting the wave peaks, relatively accurate vehicle position points can be obtained. Fig. 4 displays the raw optical fiber signal within the first 5 minutes from the starting position of the optical fiber, along with the processing results. In the figure, the dummy line represents the original signal amplitude, the solid line represents the signal after the Butterworth filter, and the red points represent the detected peaks.
Figure 4: Raw fiber signal and filtered signal. (a) raw optical fiber signal of the starting position, (b) signal filtered by Butterworth Lowpass Filter and searched peaks.

### v2

@@ Abstract
In the realm of intelligent transportation systems, accurate and reliable traffic monitoring is crucial. Traditional devices, such as cameras and lidars, face limitations in adverse weather conditions and complex traffic scenarios, prompting the need for more resilient technologies. This paper presents traffic flow monitoring method using optical fiber-based distributed acoustic sensors (DAS). An innovative vehicle trajectory extraction algorithm is proposed to derive traffic flow statistics. In the processing of optical fiber waterfall diagrams, Butterworth low-pass filter and peaks location search algorithm are employed to determine the entry position of vehicles. Subsequently, line-by-line matching algorithm is proposed to effectively track the trajectories. Experiments were conducted in highway, tunnel and city scenarios. Visualizations show that our approach not only extracts vehicle trajectories more accurately than the classical Hough and Radon transform-based methods and MUSIC beamforming algorithm, but also facilitates the calculation of traffic flow information using the low-cost acoustic sensors. It provides a new reliable means for traffic flow monitoring which can be integrated with existing methods like vision-based method.
keywords:
[…]
1 Introduction
@@
Figure 1: Comparison of vision-based and fiber-based systems. (a) Traffic monitoring under environmental conditions that cameras constraints. (b) A schematic diagram of a multi-lane highway using optical fiber systems.
@@
Nowadays, methods for monitoring traffic flow have predominantly turned to technologies such as cameras, radars, and lidars [5]. Deep Learning algorithms, including YOLO (2D multi-object detection) [6] and CenterPoint (3D multi-object detection) [7], are increasingly utilized for vehicle detection. Additionally, an effective combination of the Kalman filter and the Hungarian algorithm can be harnessed for state estimation and data association, thereby facilitating vehicle tracking and real-time positioning [8]. Nonetheless, as shown in Fig. 1(a), the efficacy of these devices can be severely compromised by environmental conditions and deployment constraints. Compounding these issues, cameras, given their fixed viewing angles, introduce dead zones and issues relating to object sizes [9]. These problems are accentuated when detecting smaller objects nearby or larger objects at extended distances, as illustrated in Fig 1 (a). Present 2D detection methods often fall short in identifying distant small objects. While lidars can discern distance information over longer ranges, their high costs hinder widespread deployment, especially at highway ends [10]. Contrasting with cameras and lidars, which serve as “visual detectors”, underground optical fibers can be likened to “acoustic detectors”. As shown in Fig 1 (b), when a vehicle passes by, the optical fiber records its location coordinates in the form of a wavelet, which is not affected by adverse environmental conditions, and there are no dead spots. Benefiting from numerous design advantages such as resistance to electromagnetic interference, heightened sensitivity, cost-effectiveness, durability, stealth, and adaptability, optical fibers may provide another reliable way for continuous and comprehensive traffic state perception in smart highways [11] and can be integrated with camera and Lidar-based methods.
@@
In this study, we investigate optical fiber-based distributed acoustic sensing (DAS) for monitoring traffic in highway, tunnel and city scenarios. The proposed methods mainly include two parts: (1) vehicle detection and tracking by signal processing; and (2) traffic density, flow, and velocity estimation. After preprocessing optical fiber waterfall diagrams, we propose a vehicle position detection algorithm based on the Butterworth low-pass filter and peaks location search. Then vehicle trajectories are extracted based on line-by-line matching and polynomial fitting, thus achieving real-time detection and overall tracking for vehicles. We then present the calculation methods for three traffic flow information (i.e. velocity, flow and density) on road profile and road segment respectively, realizing traffic flow and speed monitoring.
@@
The performances of our proposed algorithms are verified in real datasets. We use self-established dataset for highway and tunnel scenarios, and public dataset from [12] for city scenario. The vehicle tracking results extracted by our methods are more accurate than the traditional transform-based Hough [13, 14], Radon [15] and MUSIC beamforming method [12]. In the tunnel scenario, our algorithm accuracy is up to 90.90%, which validates the usability of the proposed methods in practical scenarios considering the very low-cost signal processing-based method. Even facing complex vehicle behaviors like congestion and illegal overtaking, our proposed methods can track vehicles when lines overlap or intersect, validating the effectiveness of the proposed methods for complicated vehicle behaviors.
In this article, we explore the applicability of DAS systems in traffic monitoring and the main contributions are as follows:
@@
    To the best of our knowledge, we have pioneered the application of DAS systems in real tunnel scenario.
@@
    A novel position detection algorithm and a trajectory extortion algorithm are proposed for vehicle detection and tracking, relying solely on optical fiber signals.
@@
    Experiments in highway, tunnel and city scenarios demonstrate the effectiveness of the proposed approach, even when facing challenges like vehicle overtaking and congestion-induced crossover situations.
2 Related Works
Since DAS systems have been utilized in transportation [16], numerous exploratory efforts are emerging for traffic monitoring. In this section, we introduce theoretical analyses, existing algorithms, and application scenarios of DAS systems.
2.1 Theoretical Analysis
In our experiments, fibers were buried underground at a certain depth on one side of the road. When a vehicle passes by, the vehicle’s pressure deforms the road surface and is transmitted to the optical fiber sensor, generating a strain that can be measured for amplitude. This geodetic deformation can be depicted by the Flamant-Boussinesq model [17].
Denote a given position in three-dimensional space as $(d_{x},d_{y},d_{z})$, where $d_{x}$ and $d_{y}$ are the tangential and perpendicular distances to the road respectively, and $d_{z}$ is laying depth. The quasi-static or geodetic deformation $P(d_{x},d_{y},d_{z})$ at this position is:
$P(d_{x},d_{y},d_{z})=\frac{F}{4\pi G}\times p(d_{x},d_{y},d_{z}),$ (1)
where $F$ is the total force onto an infinite half-space, $G$ is the uniform shear modulus, and $p(d_{x},d_{y},d_{z})=\frac{d_{x}}{r^{2}}(\frac{d_{z}}{r}+\frac{2\nu-1}{1+d_{z}/r})$ with $r=\sqrt{d_{x}^{2}+d_{y}^{2}+d_{z}^{2}}$ and Poisson’s ratio $\nu$.
Consequently, the impulse response $k$ of DAS systems in response to a vehicle passing by a given DAS sensor can be expressed as:
$k=\frac{1}{l}\left[P(d_{x}+\frac{l}{2},d_{y},d_{z})-P(d_{x}-\frac{l}{2},d_{y},d_{z})\right],$ (2)
where $l$ is the DAS gauge length.
(a) $d_{y}$ change
(b) $F$ change
Figure 2: Numerical simulation of the quasi-static signals. (a) Various road-perpendicular offsets, $d_{y}$, to the fiber, (b) various vehicle loads $F$.
It can be seen from Fig. 2 that theoretical amplitudes of signals decrease with the road-perpendicular offsets, $d_{y}$ and increase with the vehicle loads, $F$.
In the actual calculation process, each vehicle has four wheel positions that generate geodetic deformation. Assuming the vehicle’s front track and length as $a$ and $b$ respectively, we can then calculate the geodetic deformation at each of the four wheels. Impulse response $k$ can be defined as the difference between the response $k_{x_{1}}$ at the front end of the vehicle $d_{x_{1}}=d_{x}+\frac{l}{2}$ and the response $k_{x_{2}}$ at the rear end of the vehicle $d_{x_{2}}=d_{x}-\frac{l}{2}$:
$k=|k_{x_{2}}-k_{x_{1}}|,$ (3)
where
$k_{x_{1}}=\sum_{i=1}^{4}w_{i}\cdot P(d_{x_{1}}+\alpha_{i},d_{y}+\beta_{i},d_{z}),$ (4)
$k_{x_{2}}=\sum_{i=1}^{4}w_{i}\cdot P(d_{x_{2}}+\alpha_{i},d_{y}+\beta_{i},d_{z}).$ (5)
Here, $w_{i}$ represents the weight of the $i$-th wheel, and $(\alpha_{i},\beta_{i})$ represents the coordinates of the $i$-th wheel relative to the vehicle center. From left front to right rear of a vehicle, $(\alpha_{i},\beta_{i})$ are $(\frac{b}{2},\frac{a}{2}),(\frac{b}{2},-\frac{a}{2}),(-\frac{b}{2},-\frac{a}{2})(-\frac{b}{2},\frac{a}{2})$.
Some similar complex phenomena are also discussed in this section. We compared the simulated wave effect of a popular luxury MPV Buick GL8 ES (5.3m in length, 2.5t in load) with that of SINOTRUCK HOWO T7H (17.5m in length, 20t in load) designed for long-distance transportation, as shown in Fig. 3(a). Furthermore, we chose SINOTRUCK HOWO T7H (17.5m in length) and FAW Jiefang J6P (9.6 in length) of same weight for comparison. These two cases show similar amplitude information, which is also the problems for fiber signal acquisition to identify vehicle models.
(a) Different model with different $d_{y}$.
(b) Different length with same weight.
Figure 3: Two instances that can result in similar waveforms. It is worth noting that vehicles with longer lengths show double peaks, which is also consistent with the real situation in Fig. 7 (a) in our data set.
During actual measurements, challenges arise such as variations in optical fiber burial depths, changes in soil shear modulus, and mutual interference of vehicle signals on the road, introducing more complexities.
2.2 Existing denoising and preprocessing algorithms
Due to constraints in data acquisition, transmission, storage and other factors, fiber signals often contain significant amounts of noise, necessitating advanced denoising processing. Due to the advantages of wavelet theory in both time and frequency domains, a series of denoising algorithms are designed, such as a six-level wavelet multi-scale decomposition algorithm [18], continuous wavelet transform [19], a combination of soft and hard threshold wavelet denoising algorithm [20, 21] and wavelet threshold denoising [22]. However, these complex improved algorithms based on wavelets slow down the response speed.
In recent years, machine learning and deep learning have demonstrated significant accomplishments. Filtering algorithms and SVM are utilized for signal denoising and classification [23, 24]. An attention-based convolutional neural network architecture has been specifically devised for denoising DAS signals [25]. Time-DAE [12] and Spatial-DAE [26], based on deconvolution auto-encoder, have been proposed to deconvolve signals with impulse responses. Several kinds of novel DL methods including AlexNet, SVGG and SR-Net are designed to process DAS signals [27]. Though these algorithms can facilitate subsequent processing workflows, deep learning algorithms lack mobility and interpretability, and there are few public data sets available.
2.3 Existing vehicle detection and tracking algorithms
In the process of vehicle feature recognition and trajectory extraction, researchers often treat vehicle trajectories as straight lines and conduct domain transformations to identify the extreme points in the transformation domain. Various approaches like Hough transform [13, 14] and its improvement [28], Radon transform [15], and the MUSIC beamforming algorithm [12] are commonly employed for this purpose. However, these methods are primarily suited for extracting lines and may not be efficient for monitoring changes in vehicle positions and real-time tracking.
To address this limitation, an improved double threshold algorithm is utilized for vehicle detection [23]. Additionally, a K-means clustering algorithm and Kalman filtering technology [29] are employed for vehicle identification and tracking, with algorithm verification conducted on a simple traffic flow data subset. Furthermore, a spatial Bayesian filtering and smoothing algorithm are developed for detecting, tracking, and characterizing individual vehicles [30]. However, these algorithms tend to perform well on roads with relatively low traffic volume and uncomplicated traffic patterns.
2.4 Application scenarios in transportation
In the application scenario of DAS systems in traffic monitoring, researchers often begin with simple scenarios where vehicles pass successively, presenting relatively straightforward situations in various settings such as smart cities [31], Nanshan Iron Mine [20], villages [12], Brady Hot Springs [32], San Jose, California [30], Pasadena, California [33] during COVID-19 pandemic, as well as on campuses like Stanford campus [19] and Beijing Jiaotong University [23]. Some researchers have also explored highway scenarios, including studies like those in [25] and [13]. However, these configurations overlook challenges posed by long-distance installations and continuous tracking of vehicles. To the best of our knowledge, no studies have been conducted on DAS systems for traffic monitoring in tunnels. Besides, there have been no studies on complex vehicle behaviors such as congestion and overtaking. These scenarios present challenges within the realm of optical fiber-based traffic monitoring systems.
3 Method
In this section, we formulate the problem and propose a vehicle position detection algorithm and vehicle trajectory extraction algorithm based on optical fiber signals after preprocessing. The methodology is outlined in Fig. 4.
Figure 4: Workflow of the total algorithm. (a) Vehicle vibration data detected by DAS systems, (b) signal processing for vehicle localization and trajectory extraction, and (c) structural data for traffic monitoring.
3.1 Problem Formulation
@@
The original optical fiber data recorded by DAS systems can be represented as a matrix denoted by A:
@@
Fiber waterfall diagram: Waterfall diagrams after visualization of fiber matrix A in Eq. (6) with distance as the horizontal axis and time as the vertical axis.
@@
Fiber stripe: Clear banded stripes visible to human eyes in a fiber waterfall diagram. In this paper, fiber stripes represent the driving trajectories of the vehicles.
The primary objective of this study is to validate the effectiveness of DAS systems as a novel traffic monitoring method. Additionally, accurately distinguishing vehicle trajectories and conducting traffic flow statistics under complex conditions such as congestion and overtaking poses a challenge that needs to be addressed.
3.2 Signal Preprocessing
@@
In signal processing, denoising is the primary task. After the min-max normalization, a wavelet threshold denoising algorithm combining hard and soft threshold functions is employed.
Due to its excellent time-frequency characteristics, wavelet theory has gained widespread use in signal processing and has experienced rapid development [34]. Signal energy correlates with high-amplitude wavelet coefficients, and noise energy correlates with low-amplitude wavelet coefficients. An optimal threshold can be set to identify valuable signals above it for retention or contraction, while coefficients below it are considered noise and eliminated. Then signals are reconstructed using inverse wavelet transform to obtain a denoised signal. For the selection of the wavelet threshold function, a wavelet threshold denoising algorithm combining hard and soft threshold functions is employed to denoise vehicle vibration signals. The formula is as follows:
$\overline{{\omega}}_{j,k}=\begin{cases}{\omega}_{j,k}-a\lambda,&{\omega}_{j,k}\geq\lambda;\\ 0,&\lvert{\omega}_{j,k}\rvert<\lambda;\\ {\omega}_{j,k}+a\lambda,&{\omega}_{j,k}\leq-\lambda,\end{cases}$ (7)
@@
where ${\omega}_{j,k}$ is the wavelet coefficient and $\lambda$ is the threshold. In this paper, according to the actual situation, an ideal denoised result can be obtained when $a=0.5$.
3.3 Vehicle Position Detection